How do I find the x- and y-intercepts of this equation: x+2y=18
Answer by checkley77(12844)

Y=mX+b [m=slope & b=the Y intercept.]
x+2y=18
2y=-x+18
y=-x/2+10/2
y=-x/2+5
Set Y=0 & solve for the X intercept.
0=-x/2-5
-x/2=5
-x=5*-2
-x=-10
x=10 or (10,0)
The Y intercept=(0,5)